### v3

# Toward Clinically Grounded Foundation Models in Pathology

Abstract
In non-medical domains, foundation models (FMs) have revolutionized computer vision and language processing through large-scale self-supervised and multimodal learning. Consequently, their rapid adoption in computational pathology was expected to deliver comparable breakthroughs in cancer diagnosis, prognostication, and multimodal retrieval. However, recent systematic evaluations reveal fundamental weaknesses—low diagnostic accuracy, poor robustness, geometric instability, heavy computational demands, and concerning safety vulnerabilities. This short paper examines these shortcomings and argues that they stem from deeper conceptual mismatches between the assumptions underlying generic foundation modeling in mainstream AI and the intrinsic complexity of human tissue. Seven interrelated causes are identified: biological complexity, ineffective self-supervision, overgeneralization, excessive architectural complexity, lack of domain-specific innovation, insufficient data, and a fundamental design flaw related to tissue patch size. These findings suggest that current pathology foundation models remain conceptually misaligned with the nature of tissue morphology and call for a fundamental rethinking of the paradigm itself.
1. Introduction
Foundation models (FMs) have transformed image and text processing by leveraging massive data corpora and self-supervised learning [1, 2, 3]. Their success in general computer vision inspired enthusiasm for analogous applications in histopathology, where the potential for automating diagnostic reasoning and multimodal retrieval is expected to be immense [4, 5]. However, emerging evidence indicates that pathology FMs underperform relative to expectations and exhibit critical weaknesses under realistic evaluation. This short paper analyzes these limitations and investigates the conceptual factors contributing to their shortcomings, with the goal of informing future improvements in model design.
2. Empirical Evidence of Weaknesses
While not exhaustive, the following summaries highlight several recent studies that have reported weaknesses in foundation models to analyze tissue images.
Low Accuracy and Inconsistent Clinical Performance – Alfasly et al. [6] evaluated several leading pathology foundation models—including UNI, GigaPath, and Virchow—on 11,444 whole-slide images from 23 organs and 117 cancer subtypes in TCGA using a zero-shot retrieval framework based on Yottixel’s patch-embedding approach. Despite their scale, these models achieved only modest performance, with macro-averaged F1 scores around 40–42% for top-5 retrieval and pronounced organ-level variability: kidneys reached up to  68% (top-1 F1) [^1] Outlier organ performance, such as the unusually high kidney F1 scores, may partly reflect genuine morphological distinctiveness—but could also indicate data leakage or hidden technical bias. In pathology FMs, these effects are common and often hard to disentangle, so all high outlier results warrant careful patient-level and site-level validation., whereas lungs dropped to  21%. Aggregating patches into single WSI-level embeddings did not improve results and sometimes degraded them, suggesting loss of spatial information. Overall, while pathology foundation models outperform older CNN baselines, their absolute accuracy remains low, revealing limited generalization across tissue types and underscoring that current FMs capture texture patterns rather than true diagnostic morphology.
Lack of Robustness and Site Bias – De Jong et al. [7] systematically evaluated the robustness and generalization of ten leading pathology foundation models across multiple institutions and datasets using a newly defined Robustness Index (RI), which quantifies whether model embeddings cluster more strongly by biological class or by medical center. The RI compares within-class versus within-center similarity (with RI>>1 indicating true biological robustness). Among all tested models, only Virchow2 achieved RI $\approx$ 1.2—meaning biological structure dominated site-specific bias—whereas all others had RI << 1 (e.g., UNI $\approx$ 0.9, Phikon-v2 $\approx$ 0.7). Embeddings from most models, therefore, grouped primarily by hospital or scanner rather than by cancer type, leading to large performance drops on unseen centers. The study concludes that current pathology foundation models remain fragile, confounded, and insufficiently domain-robust, underscoring the urgent need for cross-institutional validation and bias-resilient architectures before clinical use.
Geometric Fragility – Elphick et al. [8] investigated twelve self-supervised pathology foundation models and assess how well their latent representations remain stable when image patches are rotated. The authors apply rotations in 15° increments from 0° to 360° on patches extracted from the TCGA-KIRC dataset and compute two metrics to quantify invariance: mean mutual k-nearest neighbours (m-kNN) and mean cosine distance between embeddings of non-rotated versus rotated patches. They report that the model PathDino (a small model with less than 10M parameters [9]) achieved the highest m-kNN score of 0.85, making it the most rotation-invariant by that measure, and that Hibou‑L achieved the lowest cosine distance of 0.016, indicating the best alignment for that metric. In contrast, Virchow had the lowest m-kNN ($\approx$0.53) and Phikon 2 the highest cosine distance ($\approx$0.145), indicating much poorer invariance. Importantly, the results show statistically that models trained with explicit rotation augmentation significantly outperform those without (t = 6.91; $p<0.0001$ for m-kNN; t=8.88t=-8.88; $p<0.0001$ for cosine distance). The study thus concludes that because transformer-based architectures lack an inherent rotational inductive bias, rotation augmentation in training is a necessary design choice for pathology FMs to achieve acceptable invariance.
Resource Burden and Fragile Adaptation – Mulliqi et al. [10] conducted a large-scale study using over 100,000 prostate biopsy slides (from 7,342 patients across 15 sites in 11 countries) to compare two pathology foundation models (FMs) against a task-specific (TS) end-to-end model for prostate cancer diagnosis and Gleason grading. They found that although the FMs offered utility in data-scarce settings, when enough labeled data were available the TS model matched or even outperformed the FMs. Critically, the FMs consumed up to 35× more energy than the TS model, raising sustainability concerns. Despite the purported universality of FMs, their performance did not substantially exceed TS models in the clinically validated setting, highlighting that heavy compute demands and fine-tuning instability limit the practical superiority of FMs in pathology.
Linear Probing: Because Fine-Tuning Does Not Work – Although foundation models are often promoted for their flexibility and “emergent” adaptability to new tasks, in computational pathology their downstream use is overwhelmingly limited to linear probing—training a shallow linear classifier on frozen embeddings rather than fine-tuning the model itself. This dependency arises because most pathology FMs are too large, memory-intensive, and unstable to fine-tune on moderate-sized data sets typical of clinical research (often hundreds to a few thousand slides). Recent studies confirm that full fine-tuning frequently degrades accuracy relative to linear probing due to overfitting and catastrophic forgetting [10, 11]. Yet this pragmatic retreat stands in stark contrast to the foundational premise of the FM paradigm, as articulated in Bommasani et al. (2021) [1] and many subsequent multimodal works that promise large pretrained systems should enable zero-shot and easily fine-tuned adaptation across domains. In pathology, however, this promise collapses—most “foundations” function only as static feature extractors that must be linearly probed, a situation akin to buying a Ferrari that cannot run and then purchasing a bicycle to tow it. The contradiction highlights that current FMs are less “foundations” than frozen front-ends, exposing the gap between theoretical universality and real-world usability in medical AI.
Security and Safety Vulnerabilities – Wang et al. [12] introduced Universal and Transferable Adversarial Perturbations (UTAP), imperceptible noise patterns that collapse FM embeddings across architectures. These universal attacks threaten clinical reliability. The paper shows that visually imperceptible perturbations at modest pixel bounds can severely degrade representation quality, collapsing accuracy from $\approx\!97$% to $\approx 12$% on the attacked model and transferring across black-box models (e.g., $96.42\%\rightarrow 48.69\%$, $97.23\%\rightarrow 25.52\%$), while also harming out-of-distribution datasets. These attacks break the learned feature manifold rather than merely flipping labels, are easy to learn ($\approx 900$ patches, a few minutes on an RTX,4090), and generalize widely. Thus, the high clean-test accuracy of ViT-based patch encoders can be deceptive, leaving deployed pathology systems vulnerable and posing significant safety risks to patients.
In histopathology, these perturbations pose significant safety risks and have a dual interpretation that go beyond malicious attacks that are clearly a security risk. They have a real-world noise analogue; they approximate the small, systematic variations that arise naturally in the imaging pipeline:
$\circ$ Differences in H&E staining
$\circ$ Scanner optics and illumination variability
$\circ$ Compression artifacts
$\circ$ color normalization and rescaling
$\circ$ Slide preparation imperfections
—– dust, bubbles, section thickness, etc.
$\circ$ Downstream digital processing
—- gamma correction, color-space conversion
In this view, adversarial perturbations are not malicious but diagnostic stress tests that reveal how sensitive a model is to minor pixel-level changes that can—and do—occur in laboratory and acquisition workflows.
3. Understanding the FM Limitations
This section attempts to provide a conceptual and holistic analysis of why foundation models in pathology underperform, examining the cognitive, methodological, and epistemic assumptions underlying current AI paradigms.
Underestimating the Complexity of Human Tissue – The AI community often underestimates the semantic complexity of tissue morphology (Fig.1). A child learns to recognize dogs by age two and breeds by seven (a task at which AI performs exceptionally well) [13, 14]; A pathologist—a highly trained human adult—typically requires more than twelve years of education to distinguish cancer subtypes based on tissue morphology. Unlike natural images, tissue interpretation depends on context, scale, and clinical correlation—far beyond simple object recognition.
Figure 1: AI models can recognize dogs and even distinguish among breeds—tasks that children can perform with ease. In contrast, recognizing complex tissue patterns in pathology requires an adult with more than a decade of specialized education and training.
Ineffective Self-Supervision for Tissue Images – Most early self-supervised learning frameworks were developed and validated on single-object datasets such as ImageNet, making them less suited for complex, structure-rich images like histopathology slides, which lack discrete, well-defined objects (Fig.2). The “local-global crop” assumption fails in tissue slides, where patches contain mixed or irrelevant content. As a result, models learn stain texture instead of biological patterns, weakening generalization.
Figure 2: Self-supervised learning often rests on the implicit assumption that each image represents a single, coherent object—an assumption that fails in histopathology, where multiple heterogeneous tissue structures coexist within the same field of view.
The Myth of the Universal Model – According to the No Free Lunch theorem [15], no single model excels across all problems. Expecting one FM to generalize to all organs and cancers ignores pathology’s heterogeneity. Benchmarks reveal wide organ-dependent performance swings, underscoring the limits of universal architectures.
Architectural Excess and Occam’s Razor – Modern foundation models (FMs) often pursue ever-greater depth and parameter scaling without demonstrable performance benefits. In line with Occam’s Razor [16], progress in pathology AI may instead depend on leaner, domain-structured architectures that embody the hierarchical and contextual organization of biological tissues, thereby enhancing both interpretability and generalization.
Lack of Domain-Specific Innovation – Many pathology foundation models (FMs) are direct adaptations of general-purpose frameworks such as CLIP, DINO, or MAE, retrained or fine-tuned with pathology data. However, few incorporate domain-specific mechanisms such as magnification awareness, stain-invariant representations, or morphology-aware pretext tasks, resulting in methodological stagnation and limited engagement with the underlying biological domain [17]. Despite the unique multi-scale and heterogeneous nature of tissue images, no major advances have been introduced in model topology, input preparation, or loss formulation to explicitly tailor foundation models to the structural and semantic characteristics of histopathology.
Data Deficit and Scaling Limits – CLIP was trained on 400 million image–text pairs; no pathology dataset approaches that scale. Even the largest multi-institutional archives offer fewer than one million WSIs, fragmented and inconsistently labeled. Privacy constraints exacerbate scarcity, capping the value of foundation-scale pretraining.
[…]
Figure 3: The field of view in light microscopy is traditionally quite large—approximately 2000 × 1500 pixels—and becomes vastly larger in whole-slide images (WSIs). In contrast, most AI models operate on small image patches, typically around 224 × 224 pixels.
4. Conclusions
Foundation models have undeniably transformed vision and language, and their extension into medicine marks an important milestone in the evolution of artificial intelligence. Yet, the complexities of human tissue—and the interpretive reasoning that underpins pathology—demand more than scale alone. The limitations seen in current pathology FMs are signposts guiding the next phase of progress: from data-driven generalization toward models that integrate biological understanding, and clinical context.
Moving forward, genuine innovation will come from redefining the foundations themselves—designing systems that perceive tissue as pathologists do: multi-scale, contextual, and biologically grounded, validated within transparent and rigorous clinical frameworks. With such reorientation, foundation models can mature from computational feats into trusted instruments of discovery and diagnosis, bridging the gap between algorithmic power and medical insight.